OpenStudy (anonymous):

Please help !!! Simplify: 2(a^2b^-1)^2 X (4ab^3)^4

OpenStudy (anonymous):

4a^4b^-2*216a^4b^12=864a^8b^10

OpenStudy (anonymous):

How could 4a^ 4b be ?

OpenStudy (anonymous):

sorry that would be 2a^4b^-2

OpenStudy (anonymous):

2a^4b^-2*216a^4b^12=432a^8b^10

OpenStudy (anonymous):

Thank you can you help me with the other questions too ?

OpenStudy (anonymous):

yes...i will try.....

OpenStudy (anonymous):

Okay i dont get this one : -7(2x^2+3x-2)-(x^2+4) Can you show how you did it so that i can see how to do it ?

OpenStudy (anonymous):

-14x^2-21x+14-x^2-4 =-15x^2-21x+10

OpenStudy (anonymous):

oh btw for previous problem...4^4 is 256 not 216 so it would be 512 instead of 432

OpenStudy (anonymous):

sorry...zealindaus...

OpenStudy (anonymous):

Yeh thanks I didnt bother to check if it was right. 3-(2X-1)(3X^2+X-3)-4

OpenStudy (anonymous):

3-(6x^3+2x^2-6x-3x^2-x+3)-4 =-6x^3+x^2+7x-4

OpenStudy (anonymous):

Thanks I hope its right ill check over them later. (2x+3y)(x^2-4xy+3y^2)

OpenStudy (anonymous):

2x^3-5x^2y-6xy^2+9y^3

OpenStudy (anonymous):

3(x+4)

OpenStudy (anonymous):

3x +12

OpenStudy (anonymous):

Lol yeh woop wooop got that one right XD

OpenStudy (anonymous):

Without using a calculator and showing all ur calculations, determine between which 2 integers \[\sqrt{17}\] lay

OpenStudy (anonymous):

\[\sqrt{16} = 4\] \[\sqrt{25}=5\]

OpenStudy (anonymous):

Is this right ? 4xy + 6xyz = 2(2xy)+(3xyz) ?

OpenStudy (anonymous):

Its factorizing btw

OpenStudy (anonymous):

4xy + 6xyz = 2xy( 2+3z)

OpenStudy (anonymous):

sweet_blues u r not right...that will be 2[(2xy)+(3xyz)]

OpenStudy (anonymous):

Yeah i got it now thanks sorry people its the first time i have done factorizing and simplifying.

OpenStudy (anonymous):

you should not be sorry............

OpenStudy (anonymous):

we also make mistakes....

OpenStudy (anonymous):

Okays well i have a few more questions to do before bed time if you could help me with these itd be great. I dont think you can factorize this one : mn+3n+4m+12 ?

OpenStudy (anonymous):

n(m+3)+4(m+3)=(m+3)(n+4)

OpenStudy (anonymous):

so how was the factorisation????

OpenStudy (anonymous):

Idk my teacher just told me to do these for homework but her explenation of how to do them sucked. Says its factorizing on my work sheet ?

OpenStudy (anonymous):

haven't u understand the factorisation process?

OpenStudy (anonymous):

Last one is 3-3x^2y^2 And i kinda do but finding it hard to use them.When u guys give me the awnsers i look at the question again and work out how u got the awnser u got.

OpenStudy (anonymous):

\[3-3x^2y^2=3(1-x^2y^2)=3((1)^2-\[3(1+xy)(1-xy)\](xy)^2)\]

OpenStudy (anonymous):

Wow dont you rekon thats a bit hard for a year 7 class

OpenStudy (anonymous):

if u practice a lot u will find it easy...

OpenStudy (anonymous):

Thanks dude, there was a few of solve x problem i was having trouble with yest wanna have a look ?

OpenStudy (anonymous):

happy to notice that u r my fan...

OpenStudy (anonymous):

5x-2=2x+10 3(x+7)= -2(x-9) x^2+6x-13=14 And no problem ur pure genuis man. Im more into literature than math >.<

OpenStudy (anonymous):

ist one 3x-12=3(x-4) 2nd one 5x-3 3rd one x^2+6x-27 x^2+9x-3x-27 (x+9)(x-3)

OpenStudy (anonymous):

for the second one 3(x+7)=-2(x-9) 3x+21=-2x+18 3x+2x+21-18=0 5x+3=0

OpenStudy (anonymous):

i think u need the factorisation....

OpenStudy (anonymous):

Nah ur supose to solve x lol..

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

1 st one 5x-2=2x+10 5x-2x=10+2 3x=12 x=12/2 x=4

OpenStudy (anonymous):

2nd one 5x+3=0 5x=-3 x=-3/5

OpenStudy (anonymous):

3rd one (x+9)(x-3)=0 x=-9 or 3

OpenStudy (anonymous):

oh okay yup thanks. For this one i dont know what she wants me to write ? the degrees or what?

OpenStudy (anonymous):

1.1 a hypotenuse 1.2 b hight 1.3 c base 1.4 d right angle

OpenStudy (anonymous):

Okay Can you explain how to do this one =/

OpenStudy (anonymous):

b=sqrt(119^2-59^2)=103.34 c=sqrt(b^2+41^2)=111.18

OpenStudy (anonymous):

sqrt ?

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

Oh riiiighto Sorry its really late I am really clueless about these ones >.<

OpenStudy (anonymous):

i can give u clue but u have to do it by urself..

OpenStudy (anonymous):

gradient finding formula is \[\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

Sure just dont get angry if i get is wrong like i said its really late and it have to be in by tomarrow.

OpenStudy (anonymous):

pm = 3 - -2 -4 - -5 am i even on the right path ?

OpenStudy (anonymous):

yes ur right

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

I dont know what to do next tho =C

OpenStudy (anonymous):

ur right...the slope is 5

OpenStudy (anonymous):

So thats the awnser.. ? no = 3- -2 5 + 4

OpenStudy (anonymous):

\[\frac{3-(-2)}{-4-(-5)}=\frac51=5\]

OpenStudy (anonymous):

similarly for NO it comes 5 and for MN and PO it is 0 so MNOP is a parallelogram.